Portfolio P has 30 percent invested in Stock X and 70 percent in Stock Y. The risk-free rate of interest is 6 percent and the market risk premium is 5 percent. Portfolio P has a required return of 12 percent and Stock X has a beta of 0.75. What is the beta of Stock Y?
Answer #1

Calculation of beta of stock Y:

Expected return= Risk free rate+portfolio beta*market risk premium

12= 6+Portfolio beta*5

6= Portfolio beta*5

Portfolio beta= 1.2

Beta of Stock X*weight of Stock X+ beta of stock Y*weight of Stock Y= Portfolio beta

(0.75*0.3)+ (beta of Y*0.7)= 1.2

Beta of Y*0.7= 0.975

Beta of Y= 0.975/0.7= 1.39

Beta of stock Y= 1.39